# Certified Uncertainty for Surrogate Models of Neutron Star Equations of State via Mondrian Conformal Prediction

### v1

@@ Abstract
journal: Computer Physics Communications
Keywords: neutron stars —quation of state — machine learning — conformal prediction — uncertainty quantification
@@ 1 Introduction
The neutron star (NS) equation of state (EoS) encodes the fundamental relation between pressure and energy density in ultra-dense matter, directly governing macroscopic observables such as its mass and radius, in addition to the tidal deformability [23, 31]. Constraining the EoS from multi-messenger observations — including electromagnetic signals and gravitational waves (GW) — requires repeated solutions of the Tolman–Oppenheimer–Volkoff (TOV) equations [38, 30], in conjunction with the tidal deformability equations, across large sets of candidate EoSs. This process can be computationally expensive, particularly when coupled with Bayesian inference pipelines or large-scale parameter sweeps. These aspects provide the motivation for the development of surrogate models that emulate TOV integration while preserving physical consistency [13, 22].
@@
A central challenge for such surrogates is providing certified uncertainty quantification (UQ) that remains valid under minimal assumptions. In current NS astrophysics literature, UQ has been pursued mainly through Bayesian neural networks (BNNs) [14, 21] and Gaussian processes [37], as well as posterior sampling frameworks informed by nuclear theory and GW observations [4, 9]. While these methods are powerful, they typically depend on restrictive distributional priors, scale poorly in high-dimensional parameter spaces, and do not provide finite-sample, distribution-free guarantees.
@@
Conformal Prediction (CP) [39, 3] addresses these limitations by wrapping any black-box predictor with statistically rigorous coverage guarantees under the mild assumption of data exchangeability. CP is model-agnostic, computationally lightweight, and naturally applicable to both regression and classification tasks. Variants such as Normalized CP adapt interval sizes to local difficulty estimates, while Mondrian CP [5, 6] stratifies calibration by discrete categories, yielding tighter prediction sets without sacrificing coverage. In addition, the regression-as-classification paradigm [36] extends CP’s flexibility by discretizing target spaces into ordered bins, enabling classification-based CP methods to capture heteroscedastic and multimodal uncertainties.
@@
Recent advances position CP as a general-purpose, distribution-free calibration layer for scientific machine learning. For PDE problems, conformalized physics-informed neural networks (C-PINNs) have been introduced [32]. For operator learning, conformalized operator learners—e.g., Conformalized-DeepONet [29] and risk-controlling quantile neural operators [26], built upon the original DeepONet architecture [25]—demonstrate valid and interpretable uncertainty quantification in high-dimensional physical systems. Yet, despite these advances, CP has not been applied to neutron-star EoS surrogates, nor to multitask settings that jointly address regression of continuous observables (e.g., some observables of NS like mass, radii or tidal deformabilities) and classification of discrete EoS attributes (e.g., validity, phase-transition signatures).
@@
This gap is particularly significant because CP’s finite-sample, model-agnostic guarantees can be conditioned on physically meaningful categories, enabling uncertainty estimates to be expressed directly in physical units and focused on scientifically critical regimes. In principle, Mondrian stratification based on stellar mass or compactness could yield narrower and more informative intervals in both low- and high-mass regions, while adaptive allocation schemes for the significance level $\alpha$ [10] could prioritize coverage where observational constraints are most stringent—such as the NICER credible regions for PSR J0030++0451 and PSR J0740++6620 [27, 34, 28, 35] or the LIGO–Virgo bounds on $\Lambda_{1.4}$ [1, 2]. Integrating these ideas within a multitask EoS surrogate offers a principled route to valid coverage, improved efficiency, and physically interpretable uncertainty bands, thereby strengthening the reliability of EoS inference in the era of precision neutron-star astrophysics.
@@ 2.1 Dataset
The observational layers overlaid on our curated EoS subset are shown in Figs. 13. In the mass–radius plane (Fig. 3), we include the two independent NICER inferences for PSR J0030+0451 [27, 34] and PSR J0740+6620 [28, 35], plotting their outer credible contours as solid lines and inner credible levels as dashed. In the tidal–deformability panel (Fig. 2), restricted to $0.5\leq M/M_{\odot}\leq 2.5$, we mark two EOS-agnostic point estimates at $M=1.4\,M_{\odot}$ with asymmetric $68\%$ credible intervals: a GW-only determination, $\Lambda_{1.4}=222.9^{+420.3}_{-98.9}$, and a multimessenger determination, $\Lambda_{1.4}=265.2^{+237.9}_{-104.4}$ [19]. These values are consistent with the constraint envelope derived from GW170817 and GW190425 [1, 2], reproduced in Fig. 2 for reference. Across panels, the amber background traces visualize the diversity of the valid EoS set used throughout this work.
@@
The uniform sampling intervals adopted for each EoS parameter are listed in Table 1. Pressures are expressed in cgs units and converted to geometric units via $p_{1}^{\rm geom}=10^{p_{1,\log}}/(\mathrm{Density}\,c^{2})$. At subnuclear densities, the SLy–PP EoS is attached to ensure crustal consistency [33].
@@
Notes. $p_{1,\log_{10}}$ is the base-10 logarithm of the pressure at the first transition density $\rho_{1}$. The triplet $(\Gamma_{1},\Gamma_{2},\Gamma_{3})$ denotes the adiabatic indices governing the stiffness of each polytropic segment. The transition densities $\rho_{1}$ and $\rho_{2}$ delimit the boundaries between successive segments and are fixed to match the typical nuclear-to-core regime transition identified by Read et al. [33]. The SLy–PP Equation of State combines the microscopic Skyrme–Lyon (SLy4) model for the low-density crust with a piecewise–polytropic (PP) parameterization for the high-density core, ensuring thermodynamic continuity at $\rho_{1}$ and providing a realistic description of the outer layers of neutron stars. This composite EoS scheme has become a standard choice in neutron-star modeling because it offers a smooth and causal connection between the outer crust and the core polytropic segments, while maintaining stability and compatibility with empirical nuclear constraints [33, 15].
@@
All quantities are expressed in geometric units ($G=c=1$), which absorb constants into the definitions of mass, length, and density. In this convention, one solar mass corresponds to a length scale $GM_{\odot}/c^{2}\approx 1.47664\,\mathrm{km}$. Densities are normalized as $\rho=M_{\odot}/\mathrm{Length}^{3}$, while pressures and energies share the same units as density. This formulation eliminates cumbersome factors of $G$ and $c$ in the Tolman–Oppenheimer–Volkoff equations, improving numerical stability and interpretability [38, 30].
@@
  • 2.
    Causality: $c_{s}^{2}=\mathrm{d}P/\mathrm{d}\varepsilon\leq c^{2}$, with a small tolerance margin ($\lesssim 10$$15\%$) following Read et al. [33], to account for interpolation errors in piecewise–polytropic fits;
  • 3.
    Astrophysical viability: $M_{\max}\geq 2.0\,M_{\odot}$, consistent with heavy pulsars (e.g., PSR J0740+6620; Cromartie et al. 7, Fonseca et al. 11).
@@
EoSs violating any of the constraints are labeled as invalid. The final dataset contains 40,00040{,}000 EoSs (20,00020{,}000 valid and 20,00020{,}000 invalid), balanced by design to support both classification and regression. The valid subset is used to compute normalization statistics for the regression targets, ensuring that no information leaks from calibration or test splits. This construction mirrors the philosophy of Read et al. [33], but extends it with explicit causal filters, a balanced invalid class, and a tailored sampling of $\Lambda_{1.4}$ to facilitate surrogate training under conformal prediction.
@@ 3 Model architecture
The surrogate is trained on a consolidated archive of $\sim 4\times 10^{4}$ EoSs stored as NumPy arrays, following conventions similar to Landry and Essick [20], Fujimoto et al. [12]. Each sample contains a 6-dimensional feature vector $X=\{\log_{10}p_{1},\Gamma_{1},\Gamma_{2},\Gamma_{3},\rho_{1},\rho_{2}\}$, consistent with piecewise-polytropic descriptions [33]. Features are standardized using training-only statistics and the resulting scaler is frozen for validation, test, and downstream inference.
@@
Two supervision channels are used: (i) a binary validity label $y_{\mathrm{class}}$ obtained from TOV integration under stability, causality, and $M_{\max}$ constraints; (ii) a 4-vector regression target $y_{\mathrm{reg}}=\big(M_{\max},R_{\max},R_{1.4},\Lambda_{1.4}\big)$. A boolean mask gates the regression loss, excluding samples where any scalar is undefined (e.g., TOV failure or unphysical profiles), while classification remains active for all EoSs. Following Fujimoto et al. [12], $\Lambda_{1.4}$ is modeled in $\log$-space for numerical stability.
@@
Activation and parameterization. Hidden layers use ReLU, $f(x)=\max(0,x)$, for efficient training and expressive nonlinear mappings between microscopic EoS parameters and macroscopic observables. The six polytropic parameters $\mathbf{x}=\{p_{1,\log_{10}},\Gamma_{1},\Gamma_{2},\Gamma_{3},\rho_{1},\rho_{2}\}$ define the piecewise–polytropic EoS following Read et al. [33], where $p_{1,\log_{10}}$ is the logarithmic pressure at $\rho_{1}$, $(\Gamma_{1},\Gamma_{2},\Gamma_{3})$ control the stiffness of each segment, and $(\rho_{1},\rho_{2})$ are the transition densities separating polytropic intervals.
@@ 3.1 Conformal prediction
We quantify predictive uncertainty using split conformal prediction (CP) [39, 3], employing a training set for model fitting and a disjoint calibration set to compute empirical quantiles that guarantee finite-sample coverage under exchangeability.
@@
This framework provides certified, data-adaptive uncertainty quantification for observables that strongly differentiate between viable EoSs—most notably $\Lambda_{1.4}$, whose $R^{5}$ dependence amplifies model errors [18, 17]. The resulting intervals are directly comparable to NICER and LIGO/Virgo constraints [27, 34, 28, 35, 1, 2, 19], allowing robust confrontation of surrogate predictions with observations.
@@ 4.4 External Validation with an Independent Dataset
Figure 10: MMRR curves for four externally validated EoSs with NN=valid and TOV=valid. Solid curves correspond to full TOV integrations, while the NICER credible regions for PSR J0030 and PSR J0740 are overplotted as observational constraints Miller et al. 27, Riley et al. 34, Miller et al. 28, Riley et al. 35. All tracks intersect the allowed observational domains and reach $M_{\max}\!>\!2.0\,M_{\odot}$, supporting their physical plausibility.
@@ 4.5 Benchmarking Against Related Work
Bayesian multimessenger analyses such as Han et al. [16] report posterior credible intervals for radii and tidal deformabilities under explicit prior and likelihood models. Nonparametric or hybrid approaches, e.g. Zhou et al. [40], provide flexible posterior bounds for quantities such as $M_{\max}$. Fast neural surrogates in relativistic astrophysics and modified-gravity settings [24, 8] emphasize computational throughput but generally do not supply formally calibrated uncertainty. These works together delineate the three principal paradigms in EoS inference: full Bayesian posteriors, nonparametric or hybrid distributions, and deterministic neural accelerators.
@@ 5 Discussion and Conclusions
Limitations remain. Because the surrogate is trained on three-segment piecewise–polytropic generators in the spirit of Read et al. [33], it cannot fully capture sharp phase transitions or exotic degrees of freedom. CP guarantees hold marginally under exchangeability and do not enforce physical structure (e.g., monotonicity or causality) at the surrogate level. Moreover, the present targets are limited to scalar observables rather than full functional curves.
@@ Acknowledgments
The authors acknowledge support from the Instituto Tecnológico de Aeronáutica (ITA) for providing the research environment and computational resources that enabled this work. We also thank the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES), Brazil, for financial support. This work has been done as a part of the Project INCT-Física Nuclear e Aplicações, under No. 408419/2024-5. It is also supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under Grants No. 301779/2025-2 (M.D.), No. 401565/2023-8 (Universal - C.H.L., M.D.), No. 409736/2025-2 (Universal - C.H.L., M.D.), No. 444797/2024-6 (M.D.), and Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP) under Thematic Project No. 2024/17816-8 (C.H.L., M.D.).

### v2

@@ Abstract
Keywords:
Neutron stars, equation of state, conformal prediction, uncertainty quantification
arxiv: 2602.19363
@@ 1 Introduction
The neutron star (NS) equation of state (EoS) encodes the fundamental relation between pressure and energy density in ultra-dense matter, directly governing macroscopic observables such as its mass and radius, in addition to the tidal deformability (Lattimer and Prakash, 2012; Özel and Freire, 2016). Constraining the EoS from multi-messenger observations — including electromagnetic signals and gravitational waves (GW) — requires repeated solutions of the Tolman–Oppenheimer–Volkoff (TOV) equations (Tolman, 1939; Oppenheimer and Volkoff, 1939), in conjunction with the tidal deformability equations, across large sets of candidate EoSs. This process can be computationally expensive, particularly when coupled with Bayesian inference pipelines or large-scale parameter sweeps. These aspects provide the motivation for the development of surrogate models that emulate TOV integration while preserving physical consistency (Fujimoto et al., 2020; Landry and Essick, 2020).
@@
A central challenge for such surrogates is providing certified uncertainty quantification (UQ) that remains valid under minimal assumptions. In current NS astrophysics literature, UQ has been pursued mainly through Bayesian neural networks (BNNs) (Grigorian et al., 2018; Landry et al., 2020) and Gaussian processes (Stein et al., 2021), as well as posterior sampling frameworks informed by nuclear theory and GW observations (Annala et al., 2018; Essick et al., 2020). While these methods are powerful, they typically depend on restrictive distributional priors, scale poorly in high-dimensional parameter spaces, and do not provide finite-sample, distribution-free guarantees.
@@
Conformal Prediction (CP) (Vovk et al., 2005; Angelopoulos and Bates, 2021) addresses these limitations by wrapping any black-box predictor with statistically rigorous coverage guarantees under the mild assumption of data exchangeability. CP is model-agnostic, computationally lightweight, and naturally applicable to both regression and classification tasks. Variants such as Normalized CP adapt interval sizes to local difficulty estimates, while Mondrian CP (Boström and Johansson, 2020; Boström, 2021) stratifies calibration by discrete categories, yielding tighter prediction sets without sacrificing coverage. In addition, the regression-as-classification paradigm (Romano et al., 2020) extends CP’s flexibility by discretizing target spaces into ordered bins, enabling classification-based CP methods to capture heteroscedastic and multimodal uncertainties.
@@
Recent advances position CP as a general-purpose, distribution-free calibration layer for scientific machine learning. For PDE problems, conformalized physics-informed neural networks (C-PINNs) have been introduced (Podina et al., 2024). For operator learning, conformalized operator learners—e.g., Conformalized-DeepONet (Moya et al., 2024) and risk-controlling quantile neural operators (Ma et al., 2024), built upon the original DeepONet architecture (Lu et al., 2021)—demonstrate valid and interpretable uncertainty quantification in high-dimensional physical systems. Yet, despite these advances, CP has not been applied to neutron-star EoS surrogates, nor to multitask settings that jointly address the regression of continuous observables (e.g., some observables of NS like mass, radii, or tidal deformabilities) and the classification of discrete EoS attributes (e.g., validity, phase-transition signatures).
@@
This gap is particularly significant because CP’s finite-sample, model-agnostic guaranties can be conditioned on physically meaningful categories, enabling uncertainty estimates to be expressed directly in physical units and focused on scientifically critical regimes. In principle, Mondrian stratification based on stellar mass or compactness could yield narrower and more informative intervals in both low- and high-mass regions, while adaptive allocation schemes for the significance level $\alpha$ (Fisch et al., 2021) could prioritize coverage where observational constraints are most stringent—such as the NICER credible regions for PSR J0030++0451 and PSR J0740++6620 (Miller et al., 2019; Riley et al., 2019; Miller et al., 2021; Riley et al., 2021) or the LIGO–Virgo bounds on $\Lambda_{1.4}$ (Abbott and others, 2018, 2020). Integrating these ideas within a multitask EoS surrogate offers a principled route to valid coverage, improved efficiency, and physically interpretable uncertainty bands, thereby strengthening the reliability of EoS inference in the era of precision neutron-star astrophysics.
@@ 2.1 Dataset
The observational layers overlaid on our curated EoS subset are shown in Figs. 13. In the mass–radius plane (Fig. 3), we include the two independent NICER inferences for PSR J0030+0451 (Miller et al., 2019; Riley et al., 2019) and PSR J0740+6620 (Miller et al., 2021; Riley et al., 2021), plotting their outer credible contours as solid lines and inner credible levels as dashed. In the tidal–deformability panel (Fig. 2), restricted to $0.5\leq M/M_{\odot}\leq 2.5$, we mark two EOS-agnostic point estimates at $M=1.4\,M_{\odot}$ with asymmetric $68\%$ credible intervals: a GW-only determination, $\Lambda_{1.4}=222.9^{+420.3}_{-98.9}$, and a multimessenger determination, $\Lambda_{1.4}=265.2^{+237.9}_{-104.4}$ (Huang, 2025). These values are consistent with the constraint envelope derived from GW170817 and GW190425 (Abbott and others, 2018, 2020), reproduced in Fig. 2 for reference. Across panels, the amber background traces visualize the diversity of the valid EoS set used throughout this work.
@@
The uniform sampling intervals adopted for each EoS parameter are listed in Table 1. Pressures are expressed in cgs units and converted to geometric units via $p_{1}^{\rm geom}=10^{p_{1,\log}}/(\mathrm{Density}\,c^{2})$. At subnuclear densities, the SLy–PP EoS is attached to ensure crustal consistency (Read et al., 2009).
@@
Notes. $p_{1,\log_{10}}$ is the base-10 logarithm of the pressure at the first transition density $\rho_{1}$. The triplet $(\Gamma_{1},\Gamma_{2},\Gamma_{3})$ denotes the adiabatic indices governing the stiffness of each polytropic segment. The transition densities $\rho_{1}$ and $\rho_{2}$ delimit the boundaries between successive segments and are fixed to match the typical nuclear-to-core regime transition identified by Read et al. (2009). The SLy–PP Equation of State combines the microscopic Skyrme–Lyon (SLy4) model for the low-density crust with a piecewise–polytropic (PP) parameterization for the high-density core, ensuring thermodynamic continuity at $\rho_{1}$ and providing a realistic description of the outer layers of neutron stars. This composite EoS scheme has become a standard choice in neutron-star modeling because it offers a smooth and causal connection between the outer crust and the core polytropic segments, while maintaining stability and compatibility with empirical nuclear constraints (Read et al., 2009; Haensel and Potekhin, 2001).
@@
All quantities are expressed in geometric units ($G=c=1$), which absorb constants into the definitions of mass, length, and density. In this convention, one solar mass corresponds to a length scale $GM_{\odot}/c^{2}\approx 1.47664\,\mathrm{km}$. Densities are normalized as $\rho=M_{\odot}/\mathrm{Length}^{3}$, while pressures and energies share the same units as density. This formulation eliminates cumbersome factors of $G$ and $c$ in the Tolman–Oppenheimer–Volkoff equations, improving numerical stability and interpretability (Tolman, 1939; Oppenheimer and Volkoff, 1939).
@@
  • Causality: $c_{s}^{2}=\mathrm{d}P/\mathrm{d}\varepsilon\leq c^{2}$, with a small tolerance margin ($\lesssim 10$$15\%$) following Read et al. (2009), to account for interpolation errors in piecewise–polytropic fits;
  • Astrophysical viability: $M_{\max}\geq 2.0\,M_{\odot}$, consistent with heavy pulsars (e.g., PSR J0740+6620; Cromartie et al. 2020; Fonseca et al. 2021).
@@
EoSs violating any of the constraints are labeled as invalid. The final dataset contains 40,00040{,}000 EoSs (20,00020{,}000 valid and 20,00020{,}000 invalid), balanced by design to support both classification and regression. The valid subset is used to compute normalization statistics for the regression targets, ensuring that no information leaks from calibration or test splits. This construction mirrors the philosophy of Read et al. (2009), but extends it with explicit causal filters, a balanced invalid class, and a tailored sampling of $\Lambda_{1.4}$ to facilitate surrogate training under conformal prediction.
@@ 3 Model architecture
The surrogate is trained on a consolidated archive of $\sim 4\times 10^{4}$ EoSs stored as NumPy arrays, following conventions similar to Landry and Essick (2019); Fujimoto et al. (2023). Each sample contains a 6-dimensional feature vector $X=\{\log_{10}p_{1},\Gamma_{1},\Gamma_{2},\Gamma_{3},\rho_{1},\rho_{2}\}$, consistent with piecewise-polytropic descriptions (Read et al., 2009). Features are standardized using training-only statistics and the resulting scaler is frozen for validation, test, and downstream inference.
@@
Two supervision channels are used: (i) a binary validity label $y_{\mathrm{class}}$ obtained from TOV integration under stability, causality, and $M_{\max}$ constraints; (ii) a 4-vector regression target $y_{\mathrm{reg}}=\big(M_{\max},R_{\max},R_{1.4},\Lambda_{1.4}\big)$. A boolean mask gates the regression loss, excluding samples where any scalar is undefined (e.g., TOV failure or unphysical profiles), while classification remains active for all EoSs. Following Fujimoto et al. (2023), $\Lambda_{1.4}$ is modeled in $\log$-space for numerical stability.
@@
Activation and parameterization. Hidden layers use ReLU, $f(x)=\max(0,x)$, for efficient training and expressive nonlinear mappings between microscopic EoS parameters and macroscopic observables. The six polytropic parameters $\mathbf{x}=\{p_{1,\log_{10}},\Gamma_{1},\Gamma_{2},\Gamma_{3},\rho_{1},\rho_{2}\}$ define the piecewise–polytropic EoS following Read et al. (2009), where $p_{1,\log_{10}}$ is the logarithmic pressure at $\rho_{1}$, $(\Gamma_{1},\Gamma_{2},\Gamma_{3})$ control the stiffness of each segment, and $(\rho_{1},\rho_{2})$ are the transition densities separating polytropic intervals.
@@ 3.1 Conformal prediction
We quantify predictive uncertainty using split conformal prediction (CP) (Vovk et al., 2005; Angelopoulos and Bates, 2021), employing a training set for model fitting and a disjoint calibration set to compute empirical quantiles that guarantee finite-sample coverage under exchangeability.
@@
This framework provides certified, data-adaptive uncertainty quantification for observables that strongly differentiate between viable EoSs—most notably $\Lambda_{1.4}$, whose $R^{5}$ dependence amplifies model errors (Hinderer, 2008; Hinderer et al., 2010). The resulting intervals are directly comparable to NICER and LIGO/Virgo constraints (Miller et al., 2019; Riley et al., 2019; Miller et al., 2021; Riley et al., 2021; Abbott and others, 2018, 2020; Huang, 2025), allowing robust confrontation of surrogate predictions with observations.
@@ 4.4 External Validation with an Independent Dataset
Figure 10: MMRR curves for four externally validated EoSs with NN=valid and TOV=valid. Solid curves correspond to full TOV integrations, while the NICER credible regions for PSR J0030 and PSR J0740 are overplotted as observational constraints Miller et al. 2019; Riley et al. 2019; Miller et al. 2021; Riley et al. 2021. All tracks intersect the allowed observational domains and reach $M_{\max}\!>\!2.0\,M_{\odot}$, supporting their physical plausibility.
@@ 4.5 Benchmarking Against Related Work
Bayesian multimessenger analyses such as Han et al. (2021) report posterior credible intervals for radii and tidal deformabilities under explicit prior and likelihood models. Nonparametric or hybrid approaches, e.g. Zhou et al. (2023), provide flexible posterior bounds for quantities such as $M_{\max}$. Fast neural surrogates in relativistic astrophysics and modified-gravity settings (Liodis et al., 2024; Danarianto and others, 2025) emphasize computational throughput but generally do not supply formally calibrated uncertainty. These works together delineate the three principal paradigms in EoS inference: full Bayesian posteriors, nonparametric or hybrid distributions, and deterministic neural accelerators.
@@ 5 Discussion and Conclusions
Limitations remain. Because the surrogate is trained on three-segment piecewise–polytropic generators in the spirit of Read et al. (2009), it cannot fully capture sharp phase transitions or exotic degrees of freedom. CP guarantees hold marginally under exchangeability and do not enforce physical structure (e.g., monotonicity or causality) at the surrogate level. Moreover, the present targets are limited to scalar observables rather than full functional curves.
@@ Acknowledgments
The authors acknowledge support from the Instituto Tecnológico de Aeronáutica (ITA) for providing the research environment and computational resources that enabled this work. M.M.S.M. thanks the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES), Brazil, for financial support. This work has been done as a part of the Project INCT-Física Nuclear e Aplicações, under No. 408419/2024-5. It is also supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under Grants No. 301779/2025-2 (M.D.), No. 305327/2023-2 (C.H.L.), No. 401565/2023-8 (Universal - C.H.L., M.D.), No. 409736/2025-2 (Universal - C.H.L., M.D.), No. 444797/2024-6 (M.D.), and Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP) under Thematic Project No. 2024/17816-8 (C.H.L., M.D.).